A boy is 6years old and his sister is twice as old.when the boy is 10years old,how old his sister??
Answer by Finavon(81)

Now Boy 6; Sister 2*6= 12
Boy will be 10 in 10-6= 4 years time
In 4 years time, sister will be 12+4= 16